# The integral Chow ring of $\mathcal{R}_{2}$

### v1

@@ Abstract
In this paper we compute the integral Chow ring of the moduli stack $\mathcal{R}_{2}$ of Prym pairs of genus 2 with integral coefficients.
@@ 1 Introduction
After fixing the genus $g$ of the base curve, Prym curves have a coarse moduli space, which is denoted by $R_{g}$, and it is an étale cover of the moduli space of genus $g$ smooth curves $M_{g}$, of degree $12^{2g}-$ 1. Just as for $M_{g}$, it is natural to study the geometric properties of $R_{g}$.
@@
The integral Chow groups of the Deligne-Mumford compactification ¯$\overline{\mathcal{M}}_{2}$ were independently computed in [Lar21, DLV21] and the ones of ¯$\overline{\mathcal{M}}_{3}$ after inverting $6$ by [Per23].
On the other hand, there are fewer results for the moduli stack of Prym curves, although the Picard groups with rational coefficients can be obtained from the results in [Put12]. This paper aims to start filling this gap.
We should remark that, even though there is no ambiguity on what the stack $\mathcal{M}_{g}$ should parametrize, there are two natural choices of stacks that have $R_{g}$ as a coarse moduli space (see §2.1.1 below). We will denote them by ~$\widetilde{}_{g}$ and g, and the second one is the rigidification (in the sense of [ACV03]) of the first one along the generic $\mathbb{Z}/2\mathbb{Z}$ automorphism coming from the étale double cover. This distinction had already appeared in [Cor91, BCF03].
@@ 1.1 Results and methods
In this paper, we compute the integral Chow ring of the stack $\mathcal{R}_{2}$ (defined over an algebraically closed field of characteristic different from $2$ and $3$), using the techniques of equivariant intersection theory developed in [EG98]. The first step is to find an explicit presentation of $\mathcal{R}_{2}$.
@@
Furthermore, we use $V$ to denote the standard representation of $G$ coming from the inclusion $G\subseteq\mathrm{GL}_{2}$ as the subgroup preserving the set $\{(1,0),(0,1)\}\subseteq\mathbb{C}^{2}$.
@@
where $\Delta$ is the locus of polynomials having either a root at 00 or $\infty$, or having a double root.
@@
Let $C$ be a smooth genus $2$ curve defined over an algebraically closed field $k$ of characteristic different from 2222111 In [Ver11] the field $k$ is assumed to be the field of complex numbers, but the result we are interested in remains true (with the same proof) over every algebraically closed field of characteristic different from $2$. and let $q:C\to\mathbb{P}^{1}$ be its associated double cover (well-defined up to an isomorphism of $\mathbb{P}^{1}$). Then the ramification divisor of $u$ is the set
@@
of the Weierstrass points of $C$. Let $H=q^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)$ be the hyperelliptic line bundle of $C$ and $E$ be the family of effective divisors of cardinality $2$, where which are supported on $W$. For $e\in E$ the line bundle $H\otimes\mathcal{O}_{C}(-e)$ is a non-trivial square root of $\mathcal{O}_{C}$, i.e. an element of $\mathrm{Pic}^{0}(C)[2]\smallsetminus\{\mathcal{O}_{C}\}$. The following lemma is proven in [Ver11].
@@
given by $XY\in H^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(2)\otimes\mathcal{O}_{\mathbb{P}^{1}}(-[0]-[\infty]))$.
@@
Over any base field of characteristic distinct from $2$ and $3$, the Chow ring of the moduli space of Prym pairs of genus 2 is given by
@@ Acknowledgments
We also would like to thank Younghan Bae, Samir Canning, Andrea di Lorenzo, Andrew Kresch, Sam Molcho, Michele Pernice and Johannes Schmitt for several discussions about generalities on stacks and their Chow rings, stacks of Jacobians and rigidifications. Finally, we are grateful to Gavril Farkas for conversations about the ambiguity in what the stack of Prym pairs can refer to, and to Rahul Pandharipande for inviting us to speak about this project at the ETH Algebraic Geometry and Moduli Seminars. Section 2.1.1 was added in response to his questions. A.C. and A.I.L. received support from SNF-200020-182181 and SNF-200020-219369.
@@ 2.1 Definitions
A Prym curve of genus $g$ is the datum of $(C,\eta,\beta)$ where $C$ is a smooth geometrically connected genus $g$ curve, $\eta\in\mathrm{Pic}(C)$ is non-trivial and $\beta:\eta^{\otimes 2}\to\mathcal{O}_{C}$ is an isomophism of invertible sheaves on $C$.
@@
[…]
such that there exists a isomorphism $\tau:\varphi^{*}\eta^{\prime}\to\eta$  [^2] Observe that we are adopting the convention that the datum of $\tau$ is not included in the definition of a morphism in ${}_{g}^{\mathrm{pre}}$. such that the diagram
@@ 2.1.1 Background facts about $\mathcal{R}_{g}$
Let ~$\widetilde{\mathcal{R}}_{g}$ be the stack over the big étale site whose objects are families of genus $g$ Prym curves, and a morphism $(C\to S,\eta,\beta)\to(C^{\prime}\to S^{\prime},\eta^{\prime},\beta^{\prime})$ is a Cartesian diagram of curves as in (2) and a isomorphism $\tau:\varphi^{*}\eta^{\prime}\to h^{*}\eta$ such that the diagram (3) commutes.
It is possible to show that ~$\widetilde{\mathcal{R}}_{g}$ is a smooth DM-stack. We will use this throughout.
@@
The stack ~$\widetilde{\mathcal{R}}_{g}$ has a $\mathbb{Z}/2\mathbb{Z}$-2-structure (see [AGV06, Appendix C] for the defintion) given by the multiplication by $-1$ on the line bundle $\eta$, and the corresponding rigidification is exactly $\mathcal{R}_{g}$.
Let $\widetilde{\mathcal{J}}_{g}$ be the stack of smooth curves of genus $g$ and line bundles of degree 00 with morphisms given by maps between the curves and the line bundles. Then, $\widetilde{\mathcal{J}}_{g}$ has a $\mathbb{G}_{m}$-2-structure, whose rigidification we denote by $\mathcal{J}_{g}$. The resulting stack is projective over $\mathcal{M}_{g}$ by [Del85, Theorem 4.3.].
@@
With notation as above, $\mathcal{R}_{g}$ is a smooth DM stack, with a representable, finite and étale map to $\mathcal{M}_{g}$ of degree $12^{2g}-$ 1. Moreover, there is a commutative diagram
~${\widetilde{\mathcal{R}}_{g}}$${\widetilde{\mathcal{J}}_{g}}$${\mathcal{R}_{g}}$${\mathcal{J}_{g}}$$\scriptstyle{\tilde{\iota}}$$\scriptstyle{\pi}$$\scriptstyle{\iota}$
@@
Since ~$\widetilde{\mathcal{R}}_{g}$ is a smooth DM-stack, so is $\mathcal{R}_{g}$, by [ACV03, Theorem 5.1.5.]. Moreover, the morphism $\iota:\mathcal{R}_{g}\to\mathcal{J}_{g}$ is obtained from $\tilde{\iota}$ and the universal property of rigidification [ACV03, Theorem 5.1.5.].
@@
Now, we show that it is proper. Since $\mathcal{J}_{g}\to\mathcal{M}_{g}$ is separated, it is enough to show that the natural map $\kappa:\mathcal{R}_{g}\to\mathcal{M}_{g}$ is proper. We do so using the valuative criterion for properness.
Recall that, from Remark 8, an $S$-object of $\mathcal{R}_{g}$ is a curve $C\to S$ of genus $g$, and the data $(\eta,\beta)$ of a line bundle $\eta$ and trivialization of $\eta^{\otimes 2}$ defined only after an étale base change $\widetilde{S}\to S$. Let $S=\operatorname{Spec}(R)$ be the spectrum of a DVR, and consider a smooth curve $C\to S$ of genus $g$ with the data of two Prym curves $(\eta_{i},\beta_{i})$ for $i=1,2$, defined only up to an étale cover $\widetilde{S}\to S$, that agree on the generic fiber. To check the uniqueness part of the valuative criterion, we have to show that the identity $C\to C$ is a morphism between $(C,\eta_{1},\beta_{1})$ and $(C,\eta_{2},\beta_{2})$. Because $\mathcal{J}_{g}\to\mathcal{M}_{g}$ is separated, up to a further étale cover, there is an isomorphism $\tau:\eta_{1}\to\eta_{2}$. Finally, perhaps after a further étale cover, we can also arrange $\tau$ in a way that makes the diagram (2) commute. See the proof of Proposition 25 below for more details on how to arrange $\tau$ in this way).
To check the existence part of the valuative criterion, let $C\to S$ be as before, and let $(C,\eta,\beta)$ be a $\operatorname{Frac}(R)$-point of $\mathcal{R}_{g}$. Since $\mathcal{J}_{g}$ is proper, there is an extension of DVR’s $R\to R_{1}$ such that $\eta$ is in fact a line bundle on the generic point of $C_{R_{1}}$ which can be extended to the whole $C_{R_{1}}$. Since the Picard group of a curve over a DVR and over its generic point are isomorphic, the extension is necessarily $2$-torsion too, and this shows the existence of an extension of $\beta$.
@@
Moreover, the geometric fibers of $\kappa$ have constant cardinality equal to $12^{2g}-$ 1 and, since the two stacks are smooth, this is also the degree of $\kappa$, and this implies that all fibers are reduced. Therefore $\kappa$ is finite étale.
@@
The stack ~$\widetilde{\mathcal{R}}_{g}$ is also proper over $\mathcal{M}_{g}$ (as it has the same coarse moduli space as $\mathcal{R}_{g}$), but it is not representable over $\mathcal{M}_{g}$ and does not embed into $\widetilde{\mathcal{J}}_{g}$. In fact, ~$\widetilde{\mathcal{R}}_{g}$ is DM but $\widetilde{\mathcal{J}}_{g}$ has 1-dimensional automorphism groups.
From this perspective, the stack $\mathcal{R}_{g}$ is better behaved than ~$\widetilde{\mathcal{R}}_{g}$. Both variants already appeared in the literature. The paper [Cor91] discusses ~$\widetilde{\mathcal{R}}_{g}$, while [Cor89, BCF03] focus on $\mathcal{R}_{g}$.
@@
There is nothing special about $2$-torsion in this argument. For any $n$ that is invertible over the base field, if one studies curves with a line bundle of order $n$, there are two stacks $\mathcal{M}_{g}(n)$ and ~$\widetilde{\mathcal{M}}_{g}(n)$. The first one is the $\mathbb{Z}/n\mathbb{Z}$-rigidification of the latter, and embeds into $\mathcal{J}_{g}$.
@@ 2.2 Preliminaries on $G$
We start with some notation.
Notation 14.
@@
The next proposition explain how the group $G$ arises in our computation.
@@
of automorphism of $(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(-3))$ that preserve the set $\{0,\infty\}\subseteq\mathbb{P}^{1}$.
@@
$\underline{\mathrm{Aut}}_{\{0,\infty\}}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(-1))\subset\underline{\mathrm{Aut}}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(-1))=\mathrm{GL}_{2}.$
@@
where $V_{G}$, $\Gamma_{G}$ are the vector bundles over $BG$ associated to the representations $V$ and $\Gamma$ defined in Definition 1.
@@ 2.3 Construction of the morphism
where $XY:S\to\mathrm{Sym}^{2}(V^{\vee})$ is the constant function equal to $XY$. Clearly the curve $C$ comes with a map
@@
We have two sections $\sigma_{0},\sigma_{\infty}:S\to C$, corresponding to the 00 and $\infty$ section of $\mathbb{P}^{1}_{S}\to S$. As explained in [Per23, Section 1], given a section $s:S\to\mathbb{P}^{1}_{S}$ and an $\mathcal{O}_{\mathbb{P}^{1}_{S}}$-algebra $\mathcal{A}$, one has the functorial bijective map
@@
is the subset of $\mathcal{O}_{S}$-module homomorphisms $j_{0}$ such that the two maps $j_{0}^{\otimes 2}:0^{*}\mathcal{O}_{\mathbb{P}^{1}_{S}}(-6)\to\mathcal{O}_{S}$ and $XY.F:0^{*}\mathcal{O}_{\mathbb{P}^{1}_{S}}(-6)\to\mathcal{O}_{S}$ coincide. But the last morphism is clearly 00, thus we can take $j_{0}=0$. In conclusion, the section of $\sigma_{0}$ is no further data. Similarly, the section $\sigma_{\infty}$ is no further data.
@@
As explained in [Vis98, proof of Proposition 3.1], in our situation, a isomorphism $q^{*}\mathcal{O}_{\mathbb{P}^{1}_{S}}(1)\cong\omega_{C/S}$ is functorially the same data as a isomorphism
@@
in $\mathcal{X}^{\mathrm{pre}}$ is by definition the data of a morphisms $S\to S^{\prime}$ and $g:S\to G$ such that $F^{\prime}\circ f=g.F$.
@@
Here it is crucial to keep into account how $G$ acts on the space $\mathrm{Sym}^{4}(V^{\vee})\smallsetminus\Delta$. The statement is equivalent to the identity
@@
where we wrote $g(s)=(a(s),b(s);\sigma^{i}(s))\in G(S)$ with $\sigma^{i}(s)=\delta_{0}^{i}$ as explained in Notation 14. Note that in the last equality we used Equation (13). ∎
@@
To conclude the definition of $\alpha^{\mathrm{pre}}$ on morphisms we need to provide, étale locally on $S$, a morphism $\tau$ between $\varphi^{*}\eta^{\prime}$ and $\eta$ making the diagram (3) to commute. Note that, since $g$ preserves $\{0,\infty\}$, we have

### v2

@@ Abstract
In this paper we compute the integral Chow ring of the moduli stack $\mathcal{R}_{2}$ of Prym pairs of genus 2.
MSC 2020: 14C15 (primary), 14H10
@@ 1 Introduction
After fixing the genus $g$ of the base curve, the moduli stack of Prym curves has a coarse moduli space, which is denoted by $R_{g}$, and it is a finite cover of the moduli space of genus $g$ smooth curves $M_{g}$, of degree $2^{2g}-1$. Just as for $M_{g}$, it is natural to study the geometric properties of $R_{g}$.
@@
The integral Chow groups of the Deligne-Mumford compactification $\overline{\mathcal{M}}_{2}$ were independently computed in [Lar21, DLV21] and the ones of $\overline{\mathcal{M}}_{3}$ after inverting $6$ by [Per24].
On the other hand, there are fewer results for the moduli stack of Prym curves, although the Picard groups with rational coefficients can be obtained from the results in [Put12]. This paper aims to begin filling this gap by computing the integral Chow ring of the moduli stack of genus $2$ Prym pairs. It also initiates a series of works computing the integral Chow rings of hyperelliptic Prym pairs [CL25a, CL25b], with another paper forthcoming.
We should remark that, even though there is no ambiguity on what the stack $\mathcal{M}_{g}$ should parametrize, there are two natural choices of stacks that have $R_{g}$ as a coarse moduli space (see §2.1.1 below). We will denote them by $\widetilde{\mathcal{R}}_{g}$ and $\mathcal{R}_{g}$, and the second one is the rigidification (in the sense of [ACV03]) of the first one along the generic $\mathbb{Z}/2\mathbb{Z}$ automorphism coming from the étale double cover. This distinction had already appeared in [Cor91, BCF03].
@@ 1.1 Results and methods
In this paper, we compute the integral Chow ring of the stack $\mathcal{R}_{2}$ (defined over an algebraically closed field $k$ of characteristic different from $2$ and $3$), using the techniques of equivariant intersection theory developed in [EG98]. The first step is to find an explicit presentation of $\mathcal{R}_{2}$.
@@
Furthermore, we use $V$ to denote the standard representation of $G$ coming from the inclusion $G\subseteq\mathrm{GL}_{2}$ as the subgroup preserving the set of lines $\{k(1,0),k(0,1)\}\subseteq k^{2}$.
@@
where $\Delta$ is the locus of polynomials having either a root at 0 or $\infty$, or having a double root.
@@
Let $C$ be a smooth genus $2$ curve defined over an algebraically closed field $k$ of characteristic different from 22111In [Ver11] the field $k$ is assumed to be the field of complex numbers, but the result we are interested in remains true (with the same proof) over every algebraically closed field of characteristic different from $2$. and let $q:C\to\mathbb{P}^{1}$ be its associated double cover (well-defined up to an isomorphism of $\mathbb{P}^{1}$). Then the ramification divisor of $q$ is the set
@@
of the Weierstrass points of $C$. Let $H=q^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)$ be the hyperelliptic line bundle of $C$ and $E$ be the family of effective divisors of cardinality $2$, which are supported on $W$. For $e\in E$ the line bundle $H\otimes\mathcal{O}_{C}(-e)$ is a non-trivial square root of $\mathcal{O}_{C}$, i.e. an element of $\mathrm{Pic}^{0}(C)[2]\smallsetminus\{\mathcal{O}_{C}\}$. The following lemma is proven in [Ver11].
@@
given by the polynomial $XY\in H^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(2)\otimes\mathcal{O}_{\mathbb{P}^{1}}(-[0]-[\infty]))$.
@@
Over any algebraically closed base field of characteristic distinct from $2$ and $3$, the Chow ring of the moduli space of Prym pairs of genus 2 is given by
@@ Acknowledgments
We also would like to thank Younghan Bae, Samir Canning, Andrea di Lorenzo, Andrew Kresch, Alberto Landi, Sam Molcho, Michele Pernice and Johannes Schmitt for several discussions about generalities on stacks and their Chow rings, stacks of Jacobians and rigidifications. We are grateful to Gavril Farkas for conversations about the ambiguity in what the stack of Prym pairs can refer to, and to Rahul Pandharipande for inviting us to speak about this project at the ETH Algebraic Geometry and Moduli Seminars. Section 2.1.1 was added in response to his questions. Finally, we thank the referee for carefully reading the paper and for suggesting various improvements. A.C. was supported by SNF-219369 and SNF-222363, A.I.L. received support from SNF-219369.
@@ 2.1 Definitions
A Prym curve of genus $g$ is the datum of $(C,\eta,\beta)$ where $C$ is a smooth geometrically connected genus $g$ curve, $\eta\in\mathrm{Pic}(C)$ is non-trivial and $\beta:\eta^{\otimes 2}\to\mathcal{O}_{C}$ is an isomorphism of invertible sheaves on $C$.
@@
$C{C^{\prime}}$$S{S^{\prime}}$$\scriptstyle{\varphi}{\square}$$\scriptstyle{f}$ (2)
such that there exists an isomorphism $\tau:\varphi^{*}\eta^{\prime}\to\eta$  [^2] Observe that we are adopting the convention that the datum of $\tau$ is not included in the definition of a morphism in $\mathcal{R}_{g}^{\mathrm{pre}}$. such that the diagram
@@ 2.1.1 Background facts about $\mathcal{R}_{g}$
Let $\widetilde{\mathcal{R}}_{g}$ be the stack over the big étale site whose objects are families of genus $g$ Prym curves, and a morphism $(C\to S,\eta,\beta)\to(C^{\prime}\to S^{\prime},\eta^{\prime},\beta^{\prime})$ is a Cartesian diagram of curves as in (2) and an isomorphism $\tau:\varphi^{*}\eta^{\prime}\to h^{*}\eta$ such that the diagram (3) commutes.
It is possible to show that $\widetilde{\mathcal{R}}_{g}$ is a smooth DM-stack. We will use this throughout.
@@
The stack $\widetilde{\mathcal{R}}_{g}$ has a $\mathbb{Z}/2\mathbb{Z}$-2-structure (see [AGV06, Appendix C] for the defintion) given by the multiplication by 1-1 on the line bundle $\eta$, and the corresponding rigidification is exactly $\mathcal{R}_{g}$.
Let $\widetilde{\mathcal{J}}_{g}$ be the stack of smooth curves of genus $g$ and line bundles of degree 0 with morphisms given by maps between the curves and the line bundles. Then, $\widetilde{\mathcal{J}}_{g}$ has a $\mathbb{G}_{m}$-2-structure, whose rigidification we denote by $\mathcal{J}_{g}$. The resulting stack is projective over $\mathcal{M}_{g}$ by [Del85, Theorem 4.3.].
@@
With notation as above, $\mathcal{R}_{g}$ is a smooth DM stack, with a representable, finite and étale map to $\mathcal{M}_{g}$ of degree $2^{2g}-1$. Moreover, there is a commutative diagram
$\widetilde{\mathcal{R}}_{g}$$\widetilde{\mathcal{J}}_{g}$$\mathcal{R}_{g}$$\mathcal{J}_{g}$$\scriptstyle{\tilde{\iota}}$$\scriptstyle{\pi}$$\scriptstyle{\iota}$
@@
Since $\widetilde{\mathcal{R}}_{g}$ is a smooth DM-stack, so is $\mathcal{R}_{g}$, by [ACV03, Theorem 5.1.5.]. Moreover, the morphism $\iota:\mathcal{R}_{g}\to\mathcal{J}_{g}$ is obtained from $\tilde{\iota}$ and the universal property of rigidification [ACV03, Theorem 5.1.5.].
@@
Now, we show that it is proper. Recall that, from Remark 8, an $S$-object of $\mathcal{R}_{g}$ is a curve $C\to S$ of genus $g$, and the data $(\eta,\beta)$ of a line bundle $\eta$ and trivialization of $\eta^{\otimes 2}$ defined only after an étale base change $\widetilde{S}\to S$. Let $S=\operatorname{Spec}(R)$ be the spectrum of a DVR, and consider a smooth curve $C\to S$ of genus $g$ with the data of two Prym curves $(\eta_{i},\beta_{i})$ for $i=1,2$, defined only up to an étale cover $\widetilde{S}\to S$, that agree on the generic fiber. To check the uniqueness part of the valuative criterion, we have to show that the identity $C\to C$ is a morphism between $(C,\eta_{1},\beta_{1})$ and $(C,\eta_{2},\beta_{2})$. Because $\mathcal{J}_{g}\to\mathcal{M}_{g}$ is separated, up to a further étale cover, there is an isomorphism $\tau:\eta_{1}\to\eta_{2}$. Finally, perhaps after a further étale cover, we can also arrange $\tau$ in a way that makes the diagram (2) commute (see the proof of Proposition 26 below for more details on how to arrange $\tau$ in this way).
To check the existence part of the valuative criterion, let $C\to S=\mathrm{Spec}(R)$ be as before, and let $(C,\eta,\beta)$ be a $\operatorname{Frac}(R)$-point of $\mathcal{R}_{g}$. Since $\mathcal{J}_{g}$ is proper, there is an extension of DVR’s $R\to R_{1}$ such that $\eta$ is in fact a line bundle on the generic point of $C_{R_{1}}$ which can be extended to the whole $C_{R_{1}}$. Since the Picard group of a curve over a DVR and over its generic point are isomorphic, the extension is necessarily $2$-torsion too, and this shows the existence of an extension of $\beta$.
@@
Moreover, the geometric fibers of $\kappa$ have constant cardinality equal to $2^{2g}-1$ and, since the two stacks are smooth, this is also the degree of $\kappa$, and this implies that all fibers are reduced. Therefore $\kappa$ is finite étale.
@@
The stack $\widetilde{\mathcal{R}}_{g}$ is also proper over $\mathcal{M}_{g}$ (as it has the same coarse moduli space as $\mathcal{R}_{g}$), but it is not representable over $\mathcal{M}_{g}$ and does not embed into $\widetilde{\mathcal{J}}_{g}$. In fact, $\widetilde{\mathcal{R}}_{g}$ is DM but $\widetilde{\mathcal{J}}_{g}$ has 1-dimensional automorphism groups.
From this perspective, the stack $\mathcal{R}_{g}$ is better behaved than $\widetilde{\mathcal{R}}_{g}$. Both variants already appeared in the literature. The paper [Cor91] discusses $\widetilde{\mathcal{R}}_{g}$, while [Cor89, BCF03] focus on $\mathcal{R}_{g}$.
@@
There is nothing special about $2$-torsion in this argument. For any $n$ that is invertible over the base field, if one studies curves with a line bundle of order $n$, there are two stacks $\mathcal{M}_{g}(n)$ and $\widetilde{\mathcal{M}}_{g}(n)$. The first one is the $\mathbb{Z}/n\mathbb{Z}$-rigidification of the latter, and embeds into $\mathcal{J}_{g}$.
Remark 14.
As A. Landi pointed out to us, there is a canonical map $\mathcal{H}^{w}_{g,2}\to\mathcal{R}_{g}$ from the moduli stack of hyperelliptic curves with $2$-Weierstrass sections (see [EH21]), which is $\mathbb{Z}/2\mathbb{Z}$-torsor. This provides a further presentation of $\mathcal{R}_{g}$.
@@ 2.2 Preliminaries on $G$
We start with some notation. Recall that $G=(\mathbb{G}_{m}\times\mathbb{G}_{m})\rtimes\mathbb{Z}/2\mathbb{Z}$.
Notation 15.
@@
The next proposition explains how the group $G$ arises in our computation.
@@
of automorphisms of $(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(-3))$ that preserve the set $\{0,\infty\}\subseteq\mathbb{P}^{1}$.
@@
$\underline{\mathrm{Aut}}_{\{0,\infty\}}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(-1))\subset\underline{\mathrm{Aut}}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(-1))=\mathrm{GL}_{2}$
@@
where $V_{G}$, $\Gamma_{G}$ are the vector bundles over BGBG associated to the representations $V$ and $\Gamma$ defined in Definition 1.
@@ 2.3 Construction of the morphism
where $XY:S\to\mathrm{Sym}^{2}(V^{\vee})$ is the constant function equal to $XY$ and $XY.F$ denotes its product with the polynomial $F$. Clearly the curve $C$ comes with a map
@@
We have two sections $\sigma_{0},\sigma_{\infty}:S\to C$, corresponding to the 0 and $\infty$ section of $\mathbb{P}^{1}_{S}\to S$. As explained in [Per24, Section 1], given a section $s:S\to\mathbb{P}^{1}_{S}$ and an $\mathcal{O}_{\mathbb{P}^{1}_{S}}$-algebra $\mathcal{A}$, one has the functorial bijective map
@@
is the subset of $\mathcal{O}_{S}$-module homomorphisms $j_{0}$ such that the two maps $j_{0}^{\otimes 2}:0^{*}\mathcal{O}_{\mathbb{P}^{1}_{S}}(-6)\to\mathcal{O}_{S}$ and $XY.F:0^{*}\mathcal{O}_{\mathbb{P}^{1}_{S}}(-6)\to\mathcal{O}_{S}$ coincide. But the last morphism is clearly 0, thus we can take $j_{0}=0$. In conclusion, the section of $\sigma_{0}$ is no further data. Similarly, the section $\sigma_{\infty}$ is no further data.
@@
As explained in [Vis98, proof of Proposition 3.1], in our situation, an isomorphism $q^{*}\mathcal{O}_{\mathbb{P}^{1}_{S}}(1)\cong\omega_{C/S}$ is functorially the same data as an isomorphism
@@
in $\mathcal{X}^{\mathrm{pre}}$ is by definition the data of morphisms $f:S\to S^{\prime}$ and $g:S\to G$ such that $F^{\prime}\circ f=g.F$, where $g.F$ denotes the action of $g$ on $F$.
@@
Here it is crucial to take into account how $G$ acts on the space $\mathrm{Sym}^{4}(V^{\vee})\smallsetminus\Delta$. Since the algebra structure on $\mathcal{O}_{\mathbb{P}^{1}}\oplus\mathcal{O}_{\mathbb{P}^{1}}(-3)$ is determined by $XY.F$, the statement is equivalent to the identity
@@
where we wrote $g(s)=(a(s),b(s);\sigma^{i}(s))\in G(S)$ with $\sigma^{i}(s)=\delta_{0}^{i}$ as explained in Notation 15. Note that in the last equality we used Equation (13). ∎
@@
To conclude the definition of $\alpha^{\mathrm{pre}}$ on morphisms we need to provide, étale locally on $S$, a morphism $\tau$ between $\varphi^{*}\eta^{\prime}$ and $\eta$ making the diagram (3) commute. Note that, since $g$ preserves $\{0,\infty\}$, we have